# Reversible Decoupling Network for Single Image Reflection Removal

### v1

@@ Abstract
Recent deep-learning-based approaches to single-image reflection removal have shown promising advances, primarily for two reasons: 1) the utilization of recognition-pretrained features as inputs, and 2) the design of dual-stream interaction networks. However, according to the Information Bottleneck principle, high-level semantic clues tend to be compressed or discarded during layer-by-layer propagation. Additionally, interactions in dual-stream networks follow a fixed pattern across different layers, limiting overall performance. To address these limitations, we propose a novel architecture called Reversible Decoupling Network (RDNet), which employs a reversible encoder to secure valuable information while flexibly decoupling transmission- and reflection-relevant features during the forward pass. Furthermore, we customize a transmission-rate-aware prompt generator to dynamically calibrate features, further boosting performance. Extensive experiments demonstrate the superiority of RDNet over existing SOTA methods on five widely-adopted benchmark datasets. Our code will be made publicly available.
@@ 1 Introduction
Figure 1: Quantitative comparison in PSNR between ours and previous SOTA methods, where we achieve new records on all 5 datasets. Note that the scale of each axis is normalized by its second-best value. The best and second-best PSNR are displayed for reference.
Reflection is a common superimposition factor when photographing through transparent medium, such as glass. Under the circumstances, the captured image $I$ typically contains a mixture of transmission $T$ (the scene behind medium) and reflection $R$ (the reflected scene)  (Nayar et al., 1997), which can be simply expressed as $I=T+R$. The presence of reflections often hinders vital information in the transmission layer, impeding the performance of downstream computer vision tasks, such as stereo matching, optical flow, and depth estimation (Tsin et al., 2003; Yang et al., 2016; Costanzino et al., 2023). Thus, single image reflection removal/separation (SIRR) is desired to disentangle the transmission and reflection components from a single input image. However, this problem is severely ill-posed as infinitely many possible decompositions of $\hat{T}$ and $\hat{R}$ satisfy $I=\hat{T}+\hat{R}$. In other words, it is highly challenging to determine which combination is optimal if without effective priors or guidance on decomposition.
In recent years, learning-based approaches have made tremendous strides in this field (Zhang et al., 2018; Wei et al., 2019; Li et al., 2023; Hu & Guo, 2023; Zhong et al., 2024). A key consensus among these methods is to exploit hierarchical semantic representations through large-scale recognition-pretrained models, which serve as priors or regularizers during the decomposition. One pioneering deep-learning work (Zhang et al., 2018) leverages intermediate features from a pre-trained VGGNet (Simonyan & Zisserman, 2015) through the concept of hypercolumns to help differentiate between the transmission and reflection layers from mixtures. Originally from neuroscience, the term “hypercolumn” refers to a functional unit in the visual cortex that processes visual stimuli at multiple receptive-field sizes  (Hubel & Wiesel, 1974). This concept was first applied to object segmentation and localization by interpolating and stacking features extracted from different layers of a network (Hariharan et al., 2015b). However, simply mapping stacked high-dimensional hierarchies into a group of much lower-dimensional features–as input for subsequent processes–inevitably leads to considerable semantic information loss.
Previous works with SOTA performance (Hu & Guo, 2021; 2023) suggest that, all information from the source image is valuable for the task. The two components can be optimized by exchanging information between them. For any feasible decomposition ($\hat{T}$, $\hat{R}$), the following relationship holds:
@@
where $Q$ represents the information to exchange. Concretely, YTMT (Hu & Guo, 2021) and DSRNet (Hu & Guo, 2023) select $Q$ via activation functions and channel splitting, respectively. Though being effective, the information preservation is not fully guaranteed in their interaction designs, i.e., the information bottleneck induced by linear layers in YTMT and the multiplicative reductions in the gating mechanism of DSRNet.
To avoid the above risk, reversible units (Gomez et al., 2017b), which are designed to preserve information, may offer a viable solution. In particular, building coupled reversible units naturally fits the situation as follows:
$\underline{\textit{forward process}}\left\{\begin{aligned} &\hat{T}_{2}:=\hat{T}_{1}+\mathcal{F}(\hat{R}_{1})\\ &\hat{R}_{2}:=\hat{R}_{1}+\mathcal{G}(\hat{T}_{2})\\ \end{aligned}\right.;\quad\underline{\textit{reverse process}}\left\{\begin{aligned} &\hat{T}_{1}:=\hat{T}_{2}-\mathcal{F}(\hat{R}_{1})\\ &\hat{R}_{1}:=\hat{R}_{2}-\mathcal{G}(\hat{T}_{2})\\ \end{aligned}\right.,$ (2)
@@
Although the use of reversible modules can address the issue of information loss in feature interactions at the same scale, preserving multi-scale information during the feedforward process remains a challenge. Beyond the hypercolumn (Zhang et al., 2018) and the progressive hierarchy fusion (Hu & Guo, 2023), one intuitive scheme is to stack reversible modules at each scale to facilitate forward propagation while incorporating cross-scale connections to ensure effective multi-scale interaction and fusion. A straightforward approach aligning with this idea is MAXIM (Tu et al., 2022) (without consideration of information loss), which employs a fully connected mechanism across multi-scale hierarchies. Similar ideas can also be found in HRNet (Sun et al., 2019). However, operating on high-dimensional features is computationally expensive and memory-intensive.
Inspired by GLOM (Hinton, 2023), which employs a part-whole hierarchy to represent an image with multiple columns, and embodies both bottom-up and top-down interactions to mitigate the computational burden associated with fully connected layers,we integrate multi-scale feature processors into a single sub-network, referred to as a “column”. Further, we ensemble the columns in parallel and build interactions in both bottom-up and top-down manners. It is worth noting that, the scaled residual connections used in GLOM for same level interactions between adjacent columns can still cause information loss. To remedy this problem, we extend the residual connections by incorporating multi-level reversible connections, which upgrades the vanilla reversible unit (Gomez et al., 2017a).
Compared with structural designs guided by information bottleneck principle (Tishby & Zaslavsky, 2015; Hu & Guo, 2021), our proposed framework learns disentangled representations (Desjardins et al., 2012; Bengio et al., 2013) by categorizing and recombining the original information, instead of merely selecting and discarding elements, based upon a solid foundation for its information-preserving module (reversible unit). Additionally, it retains multi-scale information and facilitates cross-scale interaction. Besides, in real-world scenarios, the reflection pattern varies along with multiple factors, such as the refractive index of the transparent surface, color granularity, and viewing angle (Schechner et al., 1999). To enhance the robustness against variations in reflection strength, we further endow the model with an adaptive transmission-rate-aware prompt generator.
In light of these considerations, this paper proposes a network, called Reversible Decoupling Network (RDNet for short). The major technical contributions of this work are twofold:
@@
  • To tackle the varied reflection parameters in real-world scenarios, we introduce an adaptive transmission-rate-aware prompt generator, which learns channel scaling factors from the dataset during training and leverages this knowledge as a prior when testing. It guides the decomposition network in selecting more accurate transmission-reflection ratios, significantly enhancing the model’s generalization capabilities.
@@ 2.1 Single Image Reflection Removal
Physical formulation. In prevalent reflection removal frameworks (Levin & Weiss, 2007), an image $I$ is typically decomposed into transmission $T$ and reflection $R$ components, so as to $I=T+R$. However, in real-world scenarios, these two layers may be attenuated by factors such as diffusion and other environmental influences during superposition (Wan et al., 2020). To account for such complexities, an augmented modeling has been proposed: $I=\alpha T+\beta R$, where the coefficients $\alpha$ and $\beta$ provide adaptability to varying conditions (Wan et al., 2018; Yang et al., 2018). Nonetheless, the assumption of linear superimposition often breaks down, particularly in cases of overexposure (Wen et al., 2019). To address this concern, the concept of an alpha-matting map $W$ is incorporated, leading to a reformulation of the model as $I=W\circ T+\overline{W}\circ R$ with $\overline{W}=1-W$. While the adjustment improves the model’s flexibility, it also increases the complexity of the already ill-posed problem.
@@
The above model struggles to encapsulate the diverse reflection phenomena, highlighting the challenge of developing a universal solution. Hu and Guo (Hu & Guo, 2023) offered a more comprehensive depiction of the superimposition process by introducing a residual term: $I=\tilde{T}+\tilde{R}+\phi(T,R)$, where $\tilde{T}$ and $\tilde{R}$ signify the altered transmission and reflection information within $I$ after superimposition and degradation, as captured by camera sensors. The term $\phi(T,R)$ denotes the residual information in the reconstruction, arising from factors such as attenuation and overexposure. However, current methods primarily use the above modelings to synthesize training data, expecting the generalizability to real-world data. But, they lack explicit estimation of the physical parameters involved. Furthermore, distance-based loss functions such as mean absolute error (MAE) and mean squared error (MSE) fail to account for global color and intensity shifts. Explicitly estimating the degradation rate of the projected image could improve performance. A more detailed explanation is provided in Sec. 3.2.
Deep-learning-based modeling.
@@
Zhang et al. (Zhang et al., 2018) enhanced semantic awareness by leveraging hypercolumn features extracted from a pre-trained VGG-19 network (Hariharan et al., 2015a), together with perceptual and adversarial losses. ERRNet (Wei et al., 2019) uses misaligned pairs as training data to take a step further. But it overlooks the reflection layer, potentially increasing ambiguity in transmission recovery. Li et al. (Li et al., 2023) proposed RAGNet, a two-stage network that initially estimates the reflection component and then uses it to guide transmission prediction. Recently, the YTMT strategy proposed in (Hu & Guo, 2021) treats both components equally through a dual-stream interactive network that restores both layers simultaneously. Yet, noticing the problem hidden in the physical formulation, their interaction module relies on a linear assumption, which may upper-bound its performance. Other methods, such as BDN (Yang et al., 2018) and IBCLN (Li et al., 2020) employ reflection models with scalar weights to iteratively estimate both components, ensuring that the reflection is not too faint. However, the interaction between the two components is ignored, sometimes leading to heavy ghosting effect in transmission and reflection. Dong et al. (Dong et al., 2021) developed an iterative network that estimates a probabilistic reflection confidence map at each step. DSRNet (Hu & Guo, 2023) introduces a mutually gated interaction mechanism within a two-stage structural design. In the first stage, the network progressively fuses extracted hierarchical features, while the second stage focuses on further decomposing these features. However, the issue of information loss persists due to the multiplicative reductions in the gating mechanism. Additionally, the progressive hierarchical fusion, isolated in the first stage, does not fully ensure that the hierarchical information is preserved during the subsequent decomposition processes. Zhu et al. (Zhu et al., 2024) proposed a maximum reflection filter for estimating reflection locations and introduce a large dataset, but they similarly overlook interaction between the two layers. Our proposed RDNet addresses the drawbacks of existing approaches by incorporating reversible connections and a multi-column design.
@@ 2.2 Reversible Network
Reversible neural networks are designed to prevent information loss by enabling the recovery of original inputs from outputs, thereby maintaining data integrity. Deco and Brauer (Deco & Brauer, 1994) introduced a reversible architecture that guarantees data preservation through a residual design, which generates a lower triangular Jacobian matrix with unity diagonal elements. Building upon this concept, Dinh et al. (Dinh et al., 2015) developed the NICE framework, employing a non-linear bijective transformation between the data and a latent space. However, this design only allows volume-preserving mappings. Dinh et al. (Dinh et al., 2017) extended extended this idea by proposing a reversible transformation that does not require volume preservation. While Gomez et al. (Gomez et al., 2017a) combined the concept of invertible networks with the ResNet architecture, ensuring that each layer’s activations can be derived from the subsequent layer’s activations. This manner enables backpropagation without storing the activations in memory, except for a few non-reversible layers.
Reversible Networks for Low-level Vision. Reversible CNNs have been effectively applied to various low-level tasks, including compression (Liu et al., 2021), enhancement (Zhu et al., 2022; Wang et al., 2022; Li et al., 2022) and restoration (Huang & Dragotti, 2022; Zhu et al., 2023; Yao et al., 2023). These solutions typically employ reversible networks as a shared encoder-decoder in a generative manner, where new textures are generated to supplement ost information during degradation. However, in the task of reflection removal, the target result (the transmission image) is mixed with the reflection rather than lost. This task requires precise decoupling of the input image components instead of generating new textures. To the best of our knowledge, our work is the first to design a reversible architecture specifically for reflection removal.
@@ 3 Methodology
In this section, we present the key components of the proposed RDNet, the overall structure of which is schematically depicted in Fig. 2. Specifically, it is composed of three primary modules: the multi-column reversible encoder (MCRE), transmission-rate-aware prompt generator (TAPG) and the hierarchy decoder (HDec). The Pretrained Hierarchy Extractor (PHE) captures semantically rich hierarchical representations from the input image and transmits them to each level of the first column in MCRE. Meanwhile, TAPG learns channel-level transmission-reflection ratio priors from the data, mapping these learned fundamental parameters into prompts that guide the MCRE network. Finally, each column in MCRE employs an HDec to encode the hierarchical information, providing effective side guidance (Qin et al., 2020). The decoded hierarchies from the last column yield the final results.
@@ 3.1 Multi-scale Reversible Column Encoder
As shown in Fig. 2, our proposed Multi-Column Reversible Encoder (MCRE) employs an architecture that differs from end-to-end models (Zhang et al., 2018; Wei et al., 2019) by incorporating multiple sub-networks, each receiving column embeddings modulated by the Transmission-rate-Aware Prompt Generator (TAPG). The model is composed of a Column Embedding Layer and multiple columns that encode multi-scale information.
In MCRE, information propagation between columns is handled through two primary mechanisms: intra-level reversible connections (denoted by blue solid lines in the figure) that facilitate information preservation between columns at the same level, and inter-level connections (illustrated as red dashed lines) paired with Bidirectional Interaction Levels, enabling interactions across adjacent levels. This approach effectively decouples multi-scale features up to Level-3. As an exception, Level-4 lacks corresponding cross-level connections, conforming to the structure of the End Level. The initial column within MCRE accepts the hierarchical information extracted by the PHE, ensuring a semantic-rich representation. The subsequent multi-column reversible design ensures the lossless propagation of hierarchical information throughout the decomposition network.
Specifically, our column embedding layer employs a $7\times 7$ convolution layer with a stride of 2, producing $2\times 2$ overlapping patches $F_{-1}$ for subsequent processes. For the $i$-th column ($i\in\{1,2,...,N\}$), each level feature $F_{j}^{i},j\in\{0,1,2\}$ receives information $F_{j-1}^{i}$ from the lower level of the current column and $F_{j+1}^{i-1}$ from a higher level of the previous one. The collected features are further fused with the signal $F_{j}^{i-1}$ of the current level. The operation described above for the level $j$ is expressed as:
@@
Hierarchy Decoder. Our hierarchy decoder integrates hierarchical codes from all scales to generate the final output. We leverage several Level Decoders (LD) to interpret the higher-dimensional hierarchies with smaller resolution into lower-dimensional ones at larger resolution. The up-sampling operator in an LD is implemented by pixel-shuffle (Shi et al., 2016), an information-consistent operator before and after the scaling. The up-sampled features are then fused with the information from the previous scale with multiplication modulation. Ultimately, the final LD produces the layer residuals ($\hat{T}_{res}$ and $\hat{R}_{res}$) through another pixel-shuffle up-sampling operation and are connected with the original input to obtain the layer decomposition $\hat{T}$ and $\hat{R}$.

### v2

@@ Abstract
Recent deep-learning-based approaches to single-image reflection removal have shown promising advances, primarily for two reasons: 1) the utilization of recognition-pretrained features as inputs, and 2) the design of dual-stream interaction networks. However, according to the Information Bottleneck principle, high-level semantic clues tend to be compressed or discarded during layer-by-layer propagation. Additionally, interactions in dual-stream networks follow a fixed pattern across different layers, limiting overall performance. To address these limitations, we propose a novel architecture called Reversible Decoupling Network (RDNet), which employs a reversible encoder to secure valuable information while flexibly decoupling transmission- and reflection-relevant features during the forward pass. Furthermore, we customize a transmission-rate-aware prompt generator to dynamically calibrate features, further boosting performance. Extensive experiments demonstrate the superiority of RDNet over existing SOTA methods on five widely-adopted benchmark datasets. RDNet achieves the best performance in the NTIRE 2025 Single Image Reflection Removal in the Wild Challenge in both fidelity and perceptual comparison. Our code is available at https://github.com/lime-j/RDNet
footnotetext: †The two authors contributed equally.  *Corresponding author.
@@ 1 Introduction
Reflection is a common superimposition factor when photographing through a transparent medium like glass. In this case, the captured image $I$ contains a mixture of transmission $T$ (the scene behind medium) and reflection $R$ (the reflected scene) [28], which can be expressed as $I=T+R$. The presence of reflections often hinders vital information in the transmission layer, impeding the performance of downstream tasks, such as stereo matching, optical flow, and depth estimation [37, 45, 3, 18]. Thus, single-image reflection removal is desired to disentangle the transmission and reflection components from a single input image. However, this is severely ill-posed as infinitely many possible decompositions of $\hat{T}$ and $\hat{R}$ satisfy $I=\hat{T}+\hat{R}$. In other words, it is highly challenging to determine which combination is optimal without effective priors on decomposition.
Figure 1: Quantitative comparison in PSNR between ours and previous state-of-the-art methods, where we achieve new records on all 5 datasets. Note that the scale of each axis is normalized by its second-best value. The best and second-best PSNR values are displayed for reference.
In recent years, learning-based approaches have made tremendous strides in this field [49, 43, 24, 15, 51]. A key consensus among these methods is to exploit hierarchical semantic representations through large-scale recognition-pretrained models, which serve as priors/regularizers during training. One pioneering work [49] leverages intermediate features from a pre-trained VGGNet [34] through the concept of hypercolumns to help differentiate between the transmission and reflection layers from mixtures. Originally from neuroscience, the term “hypercolumn” refers to a functional unit in the visual cortex that processes visual stimuli at multiple receptive-field sizes  [17]. This concept was first applied to segmentation and localization by interpolating and stacking features extracted from different layers of a network [12, 50, 26]. However, simply mapping stacked high-dimensional hierarchies into a group of much lower-dimensional features–as input for subsequent processes–inevitably leads to considerable semantic information loss.
Previous works [14, 15] suggest that all information from the source image is valuable for the task. The two components can be optimized by exchanging information between them. For any feasible decomposition ($\hat{T}$, $\hat{R}$), the following relationship holds:
@@
where $Q$ represents the information to exchange, concretely, YTMT [14] and DSRNet [15] select $Q$ via activation functions and channel splitting, respectively. Though effective, the information preservation is not fully guaranteed in their interaction designs, i.e., the information bottleneck induced by linear layers in YTMT, and the multiplicative reductions in the gating mechanism of DSRNet.
To avoid the above risk, reversible units [11], which are designed to preserve information, may offer a viable solution. In particular, building coupled reversible units naturally fits the situation as follows:
$\underline{\textit{forward:}}\begin{array}[]{c}\hat{T}_{2}:=\hat{T}_{1}+\mathcal{F}(\hat{R}_{1}),\hat{R}_{2}:=\hat{R}_{1}+\mathcal{G}(\hat{T}_{2})\\ \end{array}$ (3)
$\underline{\textit{reverse:}}\begin{array}[]{c}\hat{T}_{1}:=\hat{T}_{2}-\mathcal{F}(\hat{R}_{1}),\hat{R}_{1}:=\hat{R}_{2}-\mathcal{G}(\hat{T}_{2})\\ \end{array}$ (5)
@@
Although the use of reversible modules can address the issue of information loss in feature interactions at the same scale, preserving multi-scale information during the feedforward process remains a challenge. Beyond the hypercolumn [49] and the progressive hierarchy fusion [15], one intuitive scheme is to stack reversible modules at each scale to facilitate forward propagation while incorporating cross-scale connections to ensure effective multi-scale interaction and fusion. A straightforward approach aligning with this idea is MAXIM [38] (without consideration of information loss), which employs a fully connected mechanism across multi-scale hierarchies. Similar ideas can also be found in HRNet [35]. However, operating on high-dimensional features is computationally expensive and memory-intensive.
Inspired by GLOM [13] and RevCol [2], which employ part-whole hierarchies to represent an image with multiple columns, and embody both bottom-up and top-down interactions to mitigate the computational burden associated with fully connected layers, we integrate multi-scale feature processors into a single sub-network, referred to as a “column”. Further, we ensemble the columns in parallel and build interactions in both bottom-up and top-down manners. It is worth noting that, the scaled residual connections used in GLOM for same-level interactions between adjacent columns can still cause information loss. To remedy this problem, we extend the residual connections by incorporating multi-level reversible connections [2].
Compared with structural designs guided by information bottleneck principle [36, 14], our proposed framework learns disentangled representations [6, 1] by categorizing and recombining the original information, instead of merely selecting and discarding elements, based upon a solid foundation for its information-preserving module (reversible unit). Additionally, it facilitates cross-scale interaction. Besides, in real-world scenarios, the reflection pattern varies along with multiple factors, such as the refractive index of the transparent surface, color granularity, and viewing angle [32]. To enhance the robustness against variations in reflection strength, we further endow the model with an adaptive transmission-rate-aware prompt generator.
In light of these considerations, this paper proposes a network, called Reversible Decoupling Network (RDNet for short). The technical contributions of this work are twofold:
@@
  • To tackle the varied reflection parameters in real-world scenarios, we introduce an adaptive transmission-rate-aware prompt generator, which learns channel scaling factors during training and leverages this knowledge as a prior when testing. It guides the decomposition network in selecting more accurate transmission-reflection ratios, significantly enhancing the model’s generalization capabilities.
@@ 2.1 Single Image Reflection Removal
Physical formulation. In prevalent reflection removal frameworks [20], an image $I$ is typically decomposed into transmission $T$ and reflection $R$ components, so as to $I=T+R$. However, in real-world scenarios, these two layers may be attenuated by factors such as diffusion and other environmental influences during superposition [41]. To account for such complexities, an augmented modeling has been proposed: $I=\alpha T+\beta R$, where the coefficients $\alpha$ and $\beta$ provide adaptability to varying conditions [40, 46]. Nonetheless, the assumption of linear superimposition often breaks down, particularly in cases of overexposure [44]. To address this concern, the concept of an alpha-matting map $W$ is incorporated, leading to a reformulation of the model as $I=W\circ T+\overline{W}\circ R$ with $\overline{W}=1-W$. While the adjustment improves the model’s flexibility, it also increases the complexity of the already ill-posed problem.
@@
The above model struggles to encapsulate the diverse reflection phenomena, highlighting the challenge of developing a universal solution. Hu and Guo [15] offered a more comprehensive depiction of the superimposition process by introducing a residual term: $I=\tilde{T}+\tilde{R}+\phi(T,R)$, where $\tilde{T}$ and $\tilde{R}$ signify the altered transmission and reflection information within $I$ after superimposition and degradation, as captured by camera sensors. The term $\phi(T,R)$ denotes the residual information in the reconstruction, arising from factors such as attenuation and overexposure. However, current methods primarily use the above modelings to synthesize training data, expecting the generalizability to real-world data. But, they lack explicit estimation of the physical parameters involved. Furthermore, distance-based loss functions such as mean absolute error (MAE) and mean squared error (MSE) fail to account for global color and intensity shifts. Explicitly estimating the degradation rate of the projected image could improve performance. A more detailed explanation is provided in supplementary.
@@
Deep-learning-based modeling. Zhang et al. [49] enhanced semantic awareness by leveraging hypercolumn features extracted from a pre-trained VGG-19 network [12], together with perceptual and adversarial losses. ERRNet [43] uses misaligned pairs as training data to take a step further. But it overlooks the reflection layer, potentially increasing ambiguity in transmission recovery. Li et al. [24] proposed RAGNet, a two-stage network that initially estimates the reflection component and then uses it to guide transmission prediction. Recently, the YTMT strategy proposed in [14] treats both components equally through a dual-stream interactive network that restores both layers simultaneously. Yet, noticing the problem hidden in the physical formulation, their interaction module relies on a linear assumption, which may upper-bound its performance. Other methods, such as BDN [46] and IBCLN [21] employ reflection models with scalar weights to iteratively estimate both components, ensuring that the reflection is not too faint. However, the interaction between the two components is ignored, sometimes leading to heavy ghosting effect in transmission and reflection. Dong et al. [9] developed an iterative network that estimates a probabilistic reflection confidence map at each step. DSRNet [15] introduces a mutually gated interaction mechanism within a two-stage structural design. In the first stage, the network progressively fuses extracted hierarchical features, while the second stage focuses on further decomposing these features. However, the issue of information loss persists due to the multiplicative reductions in the gating mechanism. Additionally, the progressive hierarchical fusion, isolated in the first stage, does not fully ensure that the hierarchical information is preserved during the subsequent decomposition processes. Zhu et al. [54] proposed a maximum reflection filter for estimating reflection locations and introduce a large dataset, but they similarly overlook interaction between the two layers. Our proposed RDNet addresses the drawbacks of existing approaches by incorporating reversible connections and a multi-column design.
@@ 2.2 Reversible Network
Reversible neural networks are designed to prevent information loss by enabling the recovery of original inputs from outputs, thereby maintaining data integrity. Deco and Brauer [4] introduced a reversible architecture that guarantees data preservation through a residual design, which generates a lower triangular Jacobian matrix with unity diagonal elements. Building upon this concept, Dinh et al. [7] developed the NICE framework, employing a non-linear bijective transformation between the data and a latent space. However, this design only allows volume-preserving mappings. Dinh et al. [8] extended this idea by proposing a reversible transformation that does not require volume preservation. While Gomez et al. [11] combined the concept of invertible networks with the ResNet architecture. This manner enables backpropagation without storing the activations in memory, except for a few non-reversible layers.
Reversible Networks for Low-level Vision. Reversible CNNs have been effectively applied to various low-level tasks, including compression [25], enhancement [52, 42, 23] and restoration [16, 53, 48]. These solutions typically employ reversible networks as a shared encoder-decoder in a generative manner, where new textures are generated to supplement lost information during degradation. However, in the task of reflection removal, the target result (the transmission image) is mixed with the reflection rather than lost. This task requires precise decoupling of the input image components instead of generating new textures. To the best of our knowledge, our work is the first to design a reversible architecture specifically for reflection removal.
@@ 3 Methodology
In this section, we present the key components of the proposed RDNet, the overall structure of which is schematically depicted in Fig. 2. Specifically, it is composed of three primary modules: the multi-column reversible encoder (MCRE), transmission-rate-aware prompt generator (TAPG) and the hierarchy decoder (HDec). The Pretrained Hierarchy Extractor (PHE) captures semantically rich hierarchical representations from the input image and transmits them to each level of the first column in MCRE. Meanwhile, TAPG learns channel-level transmission-reflection ratio priors from the data, mapping these learned fundamental parameters into prompts that guide the MCRE network. Finally, each column in MCRE employs an HDec to encode the hierarchical information, providing effective side guidance [30]. The decoded hierarchies from the last column yield the final results.
@@ 3.1 Multi-scale Reversible Column Encoder
As shown in Fig. 2, our proposed Multi-Column Reversible Encoder (MCRE), inspired by [2], employs an architecture that differs from end-to-end models [49, 43] by incorporating multiple sub-networks, each receiving column embeddings modulated by the Transmission-rate-Aware Prompt Generator (TAPG). The model comprises a Column Embedding Layer and multiple columns encoding multi-scale information. a mIn MCRE, information propagation between columns is handled through two primary mechanisms: intra-level reversible connections (denoted by blue solid lines in the figure) that facilitate information preservation between columns at the same level, and inter-level connections (illustrated as red dashed lines) paired with Bidirectional Interaction Levels, enabling interactions across adjacent levels. This approach effectively decouples multi-scale features up to Level-3. As an exception, Level-4 lacks corresponding cross-level connections, conforming to the structure of the End Level. The initial column within MCRE accepts the hierarchical information extracted by the PHE, ensuring a semantic-rich representation. The subsequent multi-column reversible design ensures the lossless propagation of hierarchical information throughout the decomposition network.
Specifically, our column embedding layer employs a $7\times 7$ convolution layer with a stride of 2, producing $2\times 2$ overlapping patches $F_{-1}$ for subsequent processes. Once the embedding is obtained, it is fed to each column. For the $i$-th column ($i\in\{1,2,...,N\}$), each level feature $F_{j}^{i},j\in\{0,1,2\}$ receives information $F_{j-1}^{i}$ from the lower level of the current column and $F_{j+1}^{i-1}$ from a higher level of the previous one. The collected features are further fused with the signal $F_{j}^{i-1}$ of the current level. The operation described above for the level $j$ is expressed as:
@@
(a) Input
(b) ERRNet
(c) IBCLN
(d) Dong et al.
(e) YTMT
(f) DSRNet
(g) Zhu et al.
(h) Ours
Hierarchy Decoder. Our hierarchy decoder integrates hierarchical codes from all scales to generate the final output. We leverage several Level Decoders (LD) to interpret the higher-dimensional hierarchies with smaller resolutions into lower-dimensional ones at larger resolutions. The up-sampling operator in an LD is implemented by pixel-shuffle [33], an information-consistent operator before and after the scaling. The up-sampled features are then fused with the information from the previous scale with multiplication modulation. Ultimately, the final LD produces the layer residuals ($\hat{T}_{res}$ and $\hat{R}_{res}$) through another pixel-shuffle up-sampling operation and are connected with the original input to obtain the layer decomposition $\hat{T}$ and $\hat{R}$.